Find an equation of the line that satisfies the given conditions.
y-intercept 2; parallel to the line 3x + 4y + 4 = 0

Answer by ikleyn(52937)

.
This line has an equation

3x + 4y + c = 0,     (1)


where "c" is some constant, now unknown.


To find "c", use the condition that y-intercept is 2.

It means that the point (x,y) = (0,2) lies on the line (1), i.e.

3*0 + 4*2 + c = 0,


which gives you the value of c = -8.

Answer. The equation of the straight line under the question is 3x + 4y -8 = 0.